# Interpretable Measurement of CNN Deep Feature Density using Copula and the Generalized Characteristic Function

### v2

@@ 4 Results
As we are performing copula analysis, we present separate analysis of the univariate marginal distribution, and of the multivariate copula interdependence terms. The analysis of marginals shows the distribution of feature density including the percentage of non-zero features as well as the distribution of non-zero features. The copula density interdependence term shows the interdependence of the features with the marginal distribution removed by means of a probability integral transform. The interdependence is modeled by measuring the GCF using the MOM. We present results showing qualitative description of the feature distribution as well as quantitative goodness of fit using KL-divergence and/or cross entropy loss.
@@ 4.1 Analysis of Copula Marginals
The marginals describe the univariate distribution of CNN features, and are an important component of the copula analysis. ReLU has the effect of zeroing out (deactivating) any negative-valued features such that they have no further impact on the intermediate calculations. In order to fully and adequately describe the 1D marginal distribution post-activation, we must separately model the zero-valued and non-zero features as the combination of these distributions describes the overall marginal density. The percentage of non-zero features is shown in figure 3 and also in tablular form in table 1. The percent of nonzero features varies for each architecture (resnet18, resnet50, vgg19) as well as for each dataset (imagenette2, cifar10, cifar100, and mnist). We see however a general trend amongst all models that the percent of nonzero features is quite high in layer 0, and tends to decrease in subsequent layers until layer 3 (resnet18, resnet50) or layer 2 (vgg19) before increasing slightly in subsequent layers. One primary exception to this overall shape is shown in vgg19 over imagenette2, in which the percent of nonzero features appears to decrease monotonically all the way to layer 4. The resnet models tend to start with $70-90\%$ of nonzero features in layer 0, decreasing to around $50\%$ of nonzero features by layer 4 with the exception of resnet50 mnist, which exhibits $89.6\%$ nonzero features in layer 4. The vgg19 architectures exhibit greater sparsity than the resnet18 architectures with $40-60\%$ non-zero features in layer 0 decreasing to $9-31\%$ non-zero features in layer 4. Overall these results exhibit an increase in feature sparsity (percent zeros) with network depth, although this trend is somewhat noisy often showing an uptick in nonzero percentage in the deepest layers. The magnitude and scale of these percentages are dataset dependent with vgg19 showing greater sparsity than resnet architectures.
@@
Figure 4 shows a histogram of the non-zero portion of the marginal density for resnet18 on imagenette2 for filters 0, 1, and 2. Additional plots of the non-zero marginals for resnet18, resnet50, and vgg19 are available in supplemental materials. We see in figure4 that Filters 0 (left), 1 (middle), and 2 (right) show similar shape characteristics that depend on the network depth. Importantly, we observe an interesting phenomenon, where the early layers show a more complicated univariate shape, whereas the later layers appear to more closely resemble an exponential distribution shape. The complicated shape seen in layers 0 and 1 is most likely due to contamination of the input pixel distribution into the marginal feature distribution, as these are early shallow layers in the network. In the early layers, the model is unable to transform the input pixel space very much, so we still observe remnants of this pixel distribution. However, we see in the deeper layers such as layer 3 and 4 a very clear exponential distribution of the non-zero features. We observed a similar exponential distribution consistently across other architectures and datasets with additional similar figures in supplemental materials.
Figure 4: Histogram of marginal density for pre-trained resnet18 on imagenette2 for features 0,1,2, for each of five convolutional layers. In early layers, features show some influence of original pixel distribution. Layer 0 also exhibits a few dead features (e.x. Feature 2). Deeper layers appear to exhibit an exponential distribution.
Figure 5: Quantitative goodness-of-fit of five standard distributions to the feature marginals for cifar10,cifar100,imagenette2, and mnist, across three models resnet18, resnet50, and vgg19. Shaded region shows 1-sigma confidence interval for goodness of fit. We see that exponential, gamma and Weibull are substantially and significantly better fit than uniform and Gaussian for most layers across all models and datasets, with gamma and Weibull showing the best fit but these distributions generalize the exponential distribution which shows increasingly good fit in the deeper layers of the network.
We now quantitatively compare the goodness of fit of four parametric distributions to the marginals of the non-zero features. The Distributions that we compare are the uniform distribution, the Gaussian distribution, the gamma distribution, and the Weibull distribution. The optimal parameters of these distributions are determined using the method of simulated annealing, and the goodness of fit of each of the distributions is compared using KL-divergence. Figure 5 shows the plots (with $1\sigma$ uncertainty) of each of the distributions tested over all of the architectures and datasets in our analysis. As these are $1\sigma$ confidence intervals, overlapping intervals represent an insignificant difference, whereas non-overlapping intervals represent a statistically significant difference in the goodness-of-fit. In order to prevent overfitting, the parametric distribution is fit to the training data features, whereas the KL-divergence measures the goodness of fit to the test distribution.
@@ 4.2 Analysis of Copula Interdependence
Figure 6: Select copula interdependence for pairwise features for 5 layers of resnet18 over Imagenett2. Top to bottom: Layers 0 through 4. Left: Legendre pdf. Middle: Fourier pdf. Right: Histogram pdf.
@@
Figure 6 shows the interdependence term of the copula interdependence density for resnet18 over imagenette2 for select pairwise features. Additional plots for other architectures and datasets are available in supplemental materials. Figure 6 also shows the reconstructed interdependence density function using three different methods, the GCF with normalized Legendre polynomials (left), the GCF with real-valued Fourier series (middle) and 2D histograms as a control (right). We see that all three functions appear to show very good agreement with very similar pairwise density plots for all of the feature pairs shown in figure 6.
@@
We observe that the early layers show much more complicated interdependence with greater variation, whereas later layers tend to show a similar structure for most feature pairs that we observed. Similar to our analysis of the marginals, we believe that the variation in the pairwise interdependence seen in the shallow layers (layer 0 and layer 1) is very likely due to the contamination of the input pixel and texture distribution into the shallow features. As these are early layers in the network, the network architecture is unable to fully remove the contribution of the initial pixel distribution. The later layers however show very similar pairwise interdependence. We observe a very interesting phenomenon, in which the deep features in layers 2,3 and 4 (rows 3-5 in figure 6 appear to be uncorrelated for typical values, but show a strong statistical dependence for extreme valued features. This strong statistical dependence can be observed in figure 6 (bottom three rows) by the prevalence of very high density in the upper right corner of the interdependence plot (yellow), with much lower density along the top and right edges of the plot (blue). The apparent statistical independence of the terms over typical values (green) is also clearly apparent in the pairwise interdependence plots. We also observed (shown in supplemental materials) the prevalence of deep features showing a strong anti-correlation of the extreme values, with high density (yellow) along the top left and bottom right, but low density in the upper right hand corner (blue), again with independence for the typical valued features (green). This prevalence of uncorrelated features, except in the event of extreme-valued detections, is a new observation with profound implications regarding the distribution of deep CNN features, we discuss this result further in our discussion section.
@@
Table 2: Goodness of fit of copula interdependence term for random subsets of four features on imagenette2 with resnet18. Evaluation criteria is cross entropy loss, $1\sigma$ confidence intervals are shown in parentheses. Bold shows significantly best method for each layer task.
@@
Layer 0 1.9793 (1.9587, 1.9998) 2.0036 (1.9864, 2.0207) 1.7521 (1.7209, 1.7833)
Layer 1 2.6303 (2.6299, 2.6308) 2.6290 (2.6285, 2.6295) 2.6172 (2.6164, 2.6179)
Layer 2 2.7183 (2.7182, 2.7184) 2.7192 (2.7192, 2.7193) 2.7215 (2.7214, 2.7216)
Layer 3 2.7489 (2.7486, 2.7492) 2.7435 (2.7434, 2.7435) 2.7440 (2.7439, 2.7440)
Layer 4 2.7170 (2.7170, 2.7171) 2.7192 (2.7191, 2.7192) 2.7323 (2.7322, 2.7323)
Table 3: Goodness of fit of copula interdependence term for random subsets of four features on cifar10 with resnet18. Evaluation criteria is cross entropy loss, $1\sigma$ confidence intervals are shown in parentheses. Bold shows significantly best method for each layer task.
@@
Layer 0 2.2306 (2.2174, 2.2438) 2.2359 (2.2229, 2.2490) 2.1850 (2.1712, 2.1989)
Layer 1 2.5844 (2.5830, 2.5859) 2.5812 (2.5797, 2.5827) 2.5578 (2.5548, 2.5608)
Layer 2 2.6834 (2.6828, 2.6841) 2.6852 (2.6846, 2.6858) 2.6865 (2.6858, 2.6873)
Layer 3 2.7442 (2.7440, 2.7444) 2.7461 (2.7459, 2.7463) 2.7493 (2.7491, 2.7495)
Layer 4 2.1568 (2.1502, 2.1634) 2.1268 (2.1204, 2.1332) 2.0399 (2.0339, 2.0459)
Table 4: Goodness of fit of copula interdependence term for random subsets of four features on cifar100 with resnet18. Evaluation criteria is cross entropy loss, $1\sigma$ confidence intervals are shown in parentheses. Bold shows significantly best method for each layer task.
@@
Layer 0 2.2020 (2.1812, 2.2227) 2.2085 (2.1883, 2.2288) 2.0857 (2.0644, 2.1070)
Layer 1 2.5588 (2.5571, 2.5605) 2.5600 (2.5583, 2.5618) 2.5242 (2.5219, 2.5264)
Layer 2 2.6794 (2.6788, 2.6800) 2.6806 (2.6800, 2.6812) 2.6829 (2.6822, 2.6835)
Layer 3 2.7414 (2.7414, 2.7414) 2.7415 (2.7415, 2.7415) 2.7449 (2.7449, 2.7450)
Layer 4 2.7200 (2.7198, 2.7202) 2.7220 (2.7218, 2.7222) 2.7303 (2.7301, 2.7304)
Table 5: Goodness of fit of copula interdependence term for random subsets of four features on mnist with resnet18. Evaluation criteria is cross entropy loss, $1\sigma$ confidence intervals are shown in parentheses. Bold shows significantly best method for each layer task.

### v3

@@ 4 Results
As we are performing copula analysis, we present separate analysis of the univariate marginal distribution, and of the multivariate copula interdependence terms. The analysis of marginals shows the distribution of feature density including the percentage of non-zero features, the distribution of non-zero features, as well as an analysis of the optimal Weibull tail parameter to determine if the marginals are long-tailed. The copula density interdependence term shows the interdependence of the features with the marginal distribution removed by means of a probability integral transform. The interdependence is modeled by measuring the GCF using the MOM. We present results showing qualitative description of the feature distribution as well as quantitative goodness of fit using KL-divergence and/or cross entropy loss.
@@ 4.1 Analysis of Copula Marginals
The marginals describe the univariate distribution of CNN features, and are an important component of the copula analysis. ReLU has the effect of zeroing out (deactivating) any negative-valued features such that they have no further impact on the intermediate calculations. In order to fully and adequately describe the 1D marginal distribution post-activation, we must separately model the zero-valued and non-zero features as the combination of these distributions describes the overall marginal density. The percentage of non-zero features is shown in figure 3 and also in tabular form in table 1. The percent of nonzero features varies for each architecture (ResNet-18, ResNet-50, VGG-19) as well as for each dataset (Imagenette2, CIFAR-10, CIFAR-100, and MNIST). We see however a general trend amongst all models that the percent of nonzero features is quite high in layer 0, and tends to decrease in subsequent layers until layer 3 (ResNet-18, ResNet-50) or layer 2 (VGG-19) before increasing slightly in subsequent layers. One primary exception to this overall shape is shown in VGG-19 over Imagenette2, in which the percent of nonzero features appears to decrease monotonically all the way to layer 4. The ResNet models tend to start with $70-90\%$ of nonzero features in layer 0, decreasing to around $50\%$ of nonzero features by layer 4 with the exception of ResNet-50 MNIST, which exhibits $89.6\%$ nonzero features in layer 4. The VGG-19 architectures exhibit greater sparsity than the ResNet-18 architectures with $40-60\%$ non-zero features in layer 0 decreasing to $9-31\%$ non-zero features in layer 4. Overall these results exhibit an increase in feature sparsity (percent zeros) with network depth, although this trend is somewhat noisy often showing an uptick in nonzero percentage in the deepest layers. The magnitude and scale of these percentages are dataset dependent with VGG-19 showing greater sparsity than ResNet architectures.
Figure 4 shows a histogram of the non-zero portion of the marginal density for ResNet-18 on Imagenette2 for filters 0, 1, and 2. Additional plots of the non-zero marginals for ResNet-18, ResNet-50, and VGG-19 are available in supplemental materials. We see in figure 4 that filters 0 (left), 1 (middle), and 2 (right) show similar shape characteristics that depend on the network depth. Importantly, we observe an interesting phenomenon, where the early layers show a more complicated univariate shape, whereas the later layers appear to more closely resemble an exponential distribution shape. The complicated shape seen in layers 0 and 1 is most likely due to contamination of the input pixel distribution into the marginal feature distribution, as these are early shallow layers in the network. In the early layers, the model is unable to transform the input pixel space very much, so we still observe remnants of this pixel distribution. However, we see in the deeper layers such as layer 3 and 4 a very clear exponential distribution of the non-zero features. We observed a similar exponential distribution consistently across other architectures and datasets with additional similar figures in supplemental materials.
@@
Figure 4: Histogram of marginal density for pre-trained ResNet-18 on Imagenette2 for features 0,1,2, for each of five convolutional layers. In early layers, features show some influence of original pixel distribution. Layer 0 also exhibits a few dead features (e.x. feature 2). Deeper layers appear to exhibit an exponential distribution.
Figure 5: Quantitative goodness-of-fit of five standard distributions to the feature marginals for CIFAR-10, CIFAR-100, Imagenette2, and MNIST, across three models ResNet-18, ResNet-50, and VGG-19. Shaded region shows 95% confidence interval for goodness of fit. We see that exponential, gamma and Weibull are substantially and significantly better fit than uniform and Gaussian for most layers across all models and datasets, with gamma and Weibull showing the best fit but these distributions generalize the exponential distribution which shows increasingly good fit in the deeper layers of the network.
We now quantitatively compare the goodness of fit of five parametric distributions to the marginals of the non-zero features. The distributions that we compare are the uniform distribution, the Gaussian distribution, the gamma distribution, and the Weibull distribution. The optimal parameters of these distributions are determined using the method of stochastic hill climbing, and the goodness of fit of each of the distributions is compared using KL-divergence. Figure 5 shows the plots with $95\%$ confidence intervals of the goodness-of-fit for all five distributions tested over all of the architectures and datasets in our analysis. In order to prevent overfitting, the parametric distribution is fit to the training data features, whereas the KL-divergence measures the goodness of fit to the test distribution. Additional details on the experimental procedure are available in appendix A.
@@
4.2 Analysis of The Optimal Tail Parameter
We performed a numerical analysis of the extreme value behavior of the deep feature marginals in order to determine the extent to which deep CNN features may be considered as long tailed. Our numerical analysis applies the high-threshold technique of Extreme Value Theory (EVT). Using this technique, we defined a threshold equal to the $99^{th}$ percentile of the training sample for any given feature sample. We then numerically fit the Weibull distribution with minimum 1-Wasserstein distance to the high valued sample. The details of this calculation are given in appendix B.
Figure 6: Tail parameter analysis for CIFAR-10, CIFAR-100, Imagenette2, and MNIST, across three models ResNet-18, ResNet-50, and VGG-19.
Figure 7: Empirical optimal Weibull $\theta$ tail parameters per layer (solid) versus theoretical estimates (dashed).
The theoretical analysis of Vladimirova vladimirova2019bayesian suggests that calculation of an optimal Weibull tail parameter $\theta$ as a way of determining if the marginals are indeed long-tailed. The Weibull distribution is highly flexible, because it allows for the modeling a range of tail behaviors, where $\theta\leq\frac{1}{2}$ corresponds a to a sub-Gaussian distribution, $\theta\leq 1$ corresponds to a sub-exponential distribution, and $\theta>1$ is long-tailed.
Figure 6 shows a box plot of the optimal weibull tail parameter $\theta$ layer by layer for a variety of architecture/dataset pairs. For every layer, the box and whiskers define the quantiles of the optimal Weibull $\theta$ tail parameter. We also plot a green line at $\theta=\frac{1}{2}$ and a red line at $\theta=1$. Firstly, we observe that the deepest layer (layer 4) shows a large dataset complexity dependence. For all models, the optimal $\theta$ for layer 4 is much larger for the complex datasets (CIFAR-100, Imagenette2) than it is for the simpler datasets (CIFAR-10, MNIST). This result occurs without exception in Figure 6, strongly suggesting that the distribution of sub-categories plays a role in the long-tailedness of the deepest conv layer. In literature context, Feldman and Zhang feldman2020discovering observed clearly that CIFAR-100 and ImageNet have a more prominent long-tail of rare instances, than does MNIST. One would expect that CIFAR-100 would also have a longer tail of rare instances than CIFAR-10 as is typical of datasets with a larger number of categories (zhang2025systematic, ). Nevertheless, Vladimirova vladimirova2019bayesian would suggest that feature marginals become increasingly long-tailed with depth until the very deepest layer, but this is not what we observe in Figure 6. Instead, we observe that the long-tailedness of the deepest layer depends heavily on the complexity of the dataset, with more complex datasets (CIFAR-100, Imagenette2) exhibiting longer optimal $\theta$ values than simpler datasets (CIFAR-10, MNIST).
We henceforth define the deep intermediate layers to be layers 1, 21,\;2 and $3$ for ResNet, and layers $1$ and $2$ for VGG-19. This definition is logical because, we see that layers 3 and 4 very clearly exhibit this aforementioned dataset complexity dependence for VGG-19, whereas only layer 4 exhibits this dependence for ResNet. We observe that for the median optimal $\theta$ value increases monotonically for the deep intermediate for all dataset/model pairs without exception. This strongly suggests, that the phenomenon of increasing long-tailedness with depth is at least partially true. However, the rate of increase is highly data/model dependent, and is also substantially slower than predicted by Vladimirova et al. vladimirova2019bayesian . The difference between the empirical and theoretical $\theta$ values is shown in a comparison in Figure 7. We see that the differences are indeed very large. Nevertheless, the deep intermediate do in fact show a monotonic increase in $\theta$ albeit very slowly. Notice that we define layer as after the major blocks in Figure 2, whereas Vladimirova vladimirova2019bayesian defines layer as after each convolution and/or linear layer which is much more fine grained. The plot of Figure 7 takes this difference in definition into account.
It is observed that for every model and dataset pair, at least one layer is long-tailed in that the median optimal $\theta$ parameter is greater than $1$. Moreover, all of the datasets and models with the exception of Imagenette2 ResNet-18 exhibit a long tail for the deepest intermediate layer (Layer 3 for ResNet and Layer 4 for VGG-19). For Imagenette2 ResNet-18, the deepest intermediate layer exhibits a median $\theta$ of $0.985$ which is sub-exponential. As such, at least one layer is long-tailed in all data/model pairs, and in all but one pair, the deepest intermediate layer is long-tailed.
We also observe that VGG-19 has significantly longer tails than comparable ResNet models for the deep intermediate layers. Across all datasets, all of the deep intermediate layers for VGG-19 exhibit long tails. For ResNet, only the deepest intermediate layer is long-tailed, with the exception of MNIST ResNet-18, where Layer 2 is also long-tailed, and Imagenette2 ResNet-18 where all layers are sub-exponential. As such, the deep intermediate layers of VGG-19 are noticeably longer-tailed than those of ResNet.
@@ 4.3 Analysis of Copula Interdependence
Figure 8 shows the copula interdependence term of the probability density for ResNet-18 over Imagenette2 for select pairwise features. Additional plots for other architectures and datasets are available in supplemental materials. Figure 8 also shows the reconstructed copula interdependence using three different methods, the GCF with normalized Legendre polynomials (left), the GCF with real-valued Fourier series (middle) and 2D histograms as a control (right). We see that all three functions appear to show very good agreement with very similar pairwise density plots for all of the feature pairs shown in figure 8.
@@
We observe that the early layers show much more complicated interdependence with greater variation, whereas later layers tend to show a similar structure for most feature pairs that we observed. Similar to our analysis of the marginals, we believe that the variation in the pairwise interdependence seen in the shallow layers (layer 0 and layer 1) is very likely due to the contamination of the input pixel and texture distribution into the shallow features. As these are early layers in the network, the network architecture is unable to fully remove the contribution of the initial pixel distribution. The later layers however show very similar pairwise interdependence. We observe a very interesting phenomenon, in which the deep features in layers 2,3 and 4 (rows 3-5 in figure 8) appear to be uncorrelated for typical values, but show a strong statistical dependence for extremely large valued features. This strong statistical dependence can be observed by the prevalence of very high density in the upper right corner of the interdependence plot (yellow), with much lower density along the top and right edges of the plot (blue). The apparent statistical independence of the terms over typical values (green) is also clearly seen in the pairwise interdependence plots.
In order to evaluate the goodness-of-fit of the copula independence term, we perform a supervised evaluation using cross entropy loss. Our task is to fit the copula interdependence to the training features using our method, and then we evaluate how well this describes the interdependence of the test features. For this experiment, we calculated the goodness of fit for ResNet-18 features for random groups of four features at a time for the Imagenette2, CIFAR-10, CIFAR-100 and MNIST dataset. We performed 30 rounds per experiment, and calculated the copula interdependence using 3 methods, Legendre GCF, Fourier GCF, and Histograms. The interdependence of four randomly selected features within the same layer is calculated, and compared using cross entropy loss.
For ResNet-18 on CIFAR-10, Legendre GCF works better for layers 2 and 3; however, for the latest layer, the Histogram perform best. Table 4 shows that on CIFAR-100, Legendre GCF performs well across all layers except layer 1, where Histogram outperforms. On MNIST (table 5), the Histogram method is the best for all layers except layer 3, where Legendre GCF outperforms. This is because MNIST is essentially a binary dataset with most pixels being either black or white. This binary input pixel distribution contaminates the deep feature distributions, causing a discrete quantization which can be better modeled by the discrete Histogram approach. For the other datasets (Imagenette2, CIFAR-10, CIFAR-100) with more continuous input pixel distributions, we observe in tables 2-4 that the GCF methods significantly outperform the Histogram methods for layers 2,3 and 4, with the two exceptions being layer 3 for Imagenette2 dataset, and layer 4 for the CIFAR-10 dataset. The Histogram approach is particularly sensitive to the CoD, because with $D$ dimensions and $B$ bins per dimension, the average number of samples per bin is $N/B^{D}$. Due to max-pooling, the deeper layers exhibit fewer samples relative to the shallow layers, and one would expect the GCF methods to outperform the histogram methods particularly in the situations of higher dimensionality as well as smaller sample sizes, which is consistent with our results in tables 2-4.
Figure 8: Select copula interdependence for pairwise features for 5 layers of ResNet-18 over Imagenett2. Top to bottom: Layers 0 through 4. Left: Legendre. Middle: Fourier. Right: Histogram.
@@
Table 2: Goodness of fit of copula interdependence on Imagenette2 with ResNet-18. Evaluation criteria is cross entropy loss, 95% confidence intervals are shown in parentheses.
@@
Layer 0 1.8329 (1.7882, 1.8775) 1.8854 (1.8398, 1.9309) 1.6276 (1.5710, 1.6841)
Layer 1 2.6372 (2.6353, 2.6392) 2.6376 (2.6357, 2.6395) 2.6238 (2.6213, 2.6264)
Layer 2 2.7143 (2.7140, 2.7146) 2.7146 (2.7142, 2.7149) 2.7173 (2.7170, 2.7176)
Layer 3 2.7375 (2.7374, 2.7376) 2.7367 (2.7366, 2.7369) 2.7405 (2.7403, 2.7407)
Layer 4 2.7176 (2.7174, 2.7179) 2.7197 (2.7195, 2.7199) 2.7326 (2.7324, 2.7328)
Table 3: Goodness of fit of copula interdependence term for CIFAR-10 with ResNet-18. Evaluation criteria is cross entropy loss, 95% confidence intervals are shown in parentheses.
@@
Layer 0 2.2306 (2.2048, 2.2564) 2.2359 (2.2103, 2.2616) 2.1850 (2.1580, 2.2121)
Layer 1 2.5844 (2.5816, 2.5873) 2.5812 (2.5783, 2.5841)) 2.5578 (2.5519, 2.5637)
Layer 2 2.6834 (2.6822, 2.6847) 2.6852 (2.6840, 2.6864) 2.6865 (2.6852, 2.6879)
Layer 3 2.7442 (2.7438, 2.7446) 2.7461 (2.7457, 2.7464) 2.7493 (2.7490, 2.7496)
Layer 4 2.1568 (2.1439, 2.1697) 2.1268 (2.1143, 2.1393) 2.0399 (2.0282, 2.0516))
Table 4: Goodness of fit of copula interdependence term for CIFAR-100 with ResNet-18. Evaluation criteria is cross entropy loss, 95% confidence intervals are shown in parentheses.
@@
Layer 0 2.2020 (2.1613, 2.2427) 2.2085 (2.1689, 2.2482) 2.0857 (2.0440, 2.1274)
Layer 1 2.5588 (2.5554, 2.5622) 2.5600 (2.5566, 2.5635) 2.5242 (2.5197, 2.5286)
Layer 2 2.6794 (2.6782, 2.6807) 2.6806 (2.6794, 2.6818) 2.6829 (2.6816, 2.6841)
Layer 3 2.7414 (2.7413, 2.7415) 2.7415 (2.7414, 2.7416) 2.7449 (2.7449, 2.7450)
Layer 4 2.7200 (2.7197, 2.7203) 2.7220 (2.7217, 2.7223) 2.7303 (2.7300, 2.7306)
Table 5: Goodness of fit of copula interdependence term for MNIST with ResNet-18. Evaluation criteria is cross entropy loss, 95% confidence intervals are shown in parentheses.